# Evaluating Perspectival Biases in Cross-Modal Retrieval

### v1

@@ Appendix D Example of Result from Image to Text Retrieval
CLIP-L/14 6.904 6.911 4.171 4.182 3.245 3.249 2.658 2.652
Chinese-CLIP-L/14 6.220 6.229 3.793 3.798 3.172 3.168 2.582 2.570
@@
XLM-R-L/14 0.939 0.960 0.240 0.246 0.221 0.223 0.214 0.213
XLM-R-B/16plus 0.692 0.713 0.043 0.049 0.030 0.033 0.019 0.019
@@
ColQwen2.5-3B-M 2.138 2.164 0.192 0.209 0.114 0.122 0.088 0.089
ColQwen2.5-7B-M 2.373 2.397 0.212 0.232 0.127 0.138 0.083 0.089
ColQwen2.5-v0.2 5.958 5.974 1.375 1.424 0.633 0.676 0.377 0.410
GME-Qwen2-2B 6.014 6.037 0.739 0.804 0.254 0.306 0.140 0.175
GME-Qwen2-7B 3.843 3.875 0.221 0.264 0.094 0.124 0.058 0.077
Jina-E-v4 2.859 2.886 0.157 0.186 0.072 0.093 0.045 0.059
Appendix E Language and Rank Frequency Diagram
To illustrate bias in image-to-text retrieval, we present a visualization of language groups categorized by resource level as shown in Appendix C, showing both their overall retrieval frequency as shown in Figure 13 and their frequency distribution across ranks as shown in Figure 14.
Figure 13: A language frequency histogram of each language group for all model
Figure 14: A frequency of language group at each rank for all model
Appendix F Annotator Guideline
Appendix G UMAP Analysis for Self-Preference Cultural Bias
To visualize cultural bias, the UMAP projections of text and image embeddings of all models as shown in Figure 15 and 16. The text embeddings cluster strongly by language, a proximity that supersedes semantic content. Conversely, the image embeddings do not exhibit strong country-based clustering, suggesting lower cultural bias. While other models show a similar, albeit less severe, tendency for text embeddings to be more biased than image embeddings, this effect is diminished in modern models. The GME-Qwen2 and Jina-E-v4 models only cluster very low-resource languages (Swahili, Yoruba), and the XLM-R models demonstrate superior alignment, forming a single central cluster. This discrepancy challenges retrieval systems: a query’s text embedding is biased by its language, leading the system to favor images from the same cultural context over potentially more visually relevant content from others.
Figure 16: The UMAP visualizations of the caption embeddings (left) and image embeddings (right) from the Chinese-CLIP-L/14 model applied to our dataset.
Appendix H Full Official Model Name
In this paper, we use aliases for the model names for conciseness; the full names are provided in Table 6.
Alias Used in Paper Full Model Name Parameter
CLIP-L/14 clip-vit-large-patch141 427.6M
Chinese-CLIP-L/14 Chinese-clip-vit-large-patch142 406.2M
@@ Appendix H Full Official Model Name
XLM-R-L/14 XLM-Roberta-Large-Vit-L-146 998.3M
XLM-R-L/16plus XLM-Roberta-Large-Vit-B-16Plus6 768.9M
The models are based on the following works: 1) Radford et al. [2] for CLIP-L/14; 2) Yang et al. [27] for Chinese-CLIP-L/14; 3) Faysse et al. [14] for ColQwen2 models; 4) Zhang et al. [15] for GME-Qwen2 models; 5) Günther et al. [20] for Jina-E-v4; and 6) Carlsson et al. [10] for XLM-R-VL models.
@@
Appendix I Correlation Analysis for Self-Preference Cultural Bias
We investigate how unimodal bias, which our UMAP analysis shows is more severe in the text modality as shown in Appendix G, impacts cross-modal retrieval. To quantify this, we use the Silhouette score and find that high scores in low-resource languages correlate with self-preference cultural bias score (SP) cultural bias as shown in Table 8. We confirm this relationship by calculating the Pearson correlation between SP and the Silhouette scores. For example, CLIP-L/14’s Text Silhouette score correlates strongly with SP (0.827), while its Image Silhouette correlation is only moderate (0.550), as shown in Figure 17. Across all tested models, the average correlations reveal that SP is predominantly driven by the text encoder as shown in Table 7.
Model TSC ISC
CLIP-L/14 0.83 0.55
Chinese-CLIP-L/14 -0.26 0.58
XLM-R-VL-B/16 0.98 -0.27
XLM-R-VL-L/14 0.94 0.05
Jina-E-v4 0.86 0.16
GME-Qwen2-2B 0.80 0.09
GME-Qwen2-7B 0.64 0.07
Average 0.68 0.18
Model Metrics Country
USA UK AUS GER CHN JPN FRA ESP ARG PRT BRA SAU THA IND KEN NGA
Chinese-CLIP-L/14 SP $\downarrow$ 0.03 0.06 0.05 1.11 0.03 0.13 0.33 0.40 0.43 0.56 0.52 4.88 2.55 6.98 1.99 2.23
Appendix J Dataset Statistics
@@ Appendix I Correlation Analysis for Self-Preference Cultural Bias
The distribution of cultural concepts in the XCM dataset is shown in Table 10, with each concept being represented by approximately 85 images on average.
Country Number of Samples
@@
Average 733
Table 9: Dataset Statistics of XCM dataset per Country
@@
USA UK AUS GER CHN JPN FRA ESP ARG PRT BRA SAU THA IND KEN NGA
CLIP-L/14 $M_{\text{sem}}(\%)\uparrow$ 95.73 94.57 94.73 52.42 25.17 31.07 75.53 78.00 69.65 65.78 61.33 7.75 10.48 5.56 27.83 24.19
$M_{\text{cul}}(\%)\downarrow$ 1.31 2.02 2.22 39.65 66.16 60.34 18.29 15.10 23.61 25.85 31.35 83.04 84.75 88.24 56.67 54.85
$M_{\text{non}}(\%)\downarrow$ 2.96 3.42 3.05 7.93 8.67 8.59 6.18 6.90 6.74 8.37 7.32 9.21 4.78 6.20 15.50 20.96
Chinese-CLIP -L/14 $M_{\text{sem}}(\%)\uparrow$ 94.25 89.44 90.57 40.05 93.40 84.20 65.79 61.95 60.83 56.55 57.04 14.38 21.88 10.47 27.00 25.10
$M_{\text{cul}}(\%)\downarrow$ 3.28 5.75 4.44 44.62 2.48 10.92 21.97 24.73 26.07 31.80 29.56 70.11 55.78 73.00 53.83 56.02
$M_{\text{non}}(\%)\downarrow$ 2.46 4.81 4.99 15.32 4.13 4.88 12.24 13.32 13.10 11.65 13.40 15.51 22.34 16.54 19.17 18.89
Jina-E-v4 $M_{\text{sem}}(\%)\uparrow$ 95.89 95.65 95.98 92.47 94.22 92.79 94.34 93.22 91.70 91.99 94.20 91.60 95.38 86.30 56.50 36.74
$M_{\text{cul}}(\%)\downarrow$ 1.97 2.48 1.66 4.97 3.44 3.92 4.08 3.09 4.80 4.13 3.73 4.68 2.00 9.95 27.83 34.15
$M_{\text{non}}(\%)\downarrow$ 2.13 1.86 2.36 2.55 2.34 3.29 1.58 3.69 3.50 3.88 2.07 3.72 2.62 3.75 15.67 29.11
XLM-R-L/14 $M_{\text{sem}}(\%)\uparrow$ 89.66 91.15 89.46 90.32 92.30 89.40 87.63 85.37 86.12 87.86 87.98 90.95 93.07 89.41 80.33 40.49
$M_{\text{cul}}(\%)\downarrow$ 4.11 4.04 5.83 4.44 4.13 6.26 6.58 6.54 7.39 7.16 5.80 4.85 3.54 3.75 8.50 24.71
$M_{\text{non}}(\%)\downarrow$ 6.24 4.81 4.72 5.24 3.58 4.35 5.79 8.09 6.49 4.98 6.22 4.20 3.39 6.85 11.17 34.80
XLM-R-B /16Plus $M_{\text{sem}}(\%)\uparrow$ 91.95 91.30 90.71 93.95 94.50 91.62 90.66 88.47 87.81 90.05 89.36 92.57 95.22 91.09 83.17 40.88
$M_{\text{cul}}(\%)\downarrow$ 4.11 3.88 5.13 4.03 3.58 4.56 4.34 5.47 5.97 5.70 5.52 3.88 2.00 4.91 8.33 26.78
$M_{\text{non}}(\%)\downarrow$ 3.94 4.81 4.16 2.02 1.93 3.82 5.00 6.06 6.23 4.25 5.11 3.55 2.77 4.01 8.50 32.34
GME-Qwen2 -2B-Instruct $M_{\text{sem}}(\%)\uparrow$ 96.06 95.34 96.95 87.77 94.09 93.43 92.76 90.49 88.59 90.05 88.95 84.49 89.06 76.87 37.17 25.87
$M_{\text{cul}}(\%)\downarrow$ 2.30 2.48 1.66 8.47 3.58 3.92 4.74 4.64 7.26 7.04 7.87 11.63 8.63 18.48 46.00 52.26
$M_{\text{non}}(\%)\downarrow$ 1.64 2.17 1.39 3.76 2.34 2.65 2.50 4.88 4.15 2.91 3.18 3.88 2.31 4.65 16.83 21.86
GME-Qwen2 -7B-Instruct $M_{\text{sem}}(\%)\uparrow$ 95.40 95.34 96.53 84.68 95.05 92.47 90.39 90.73 90.27 88.35 85.50 85.46 90.91 85.92 55.17 29.62
$M_{\text{cul}}(\%)\downarrow$ 2.46 2.33 0.97 11.69 3.44 4.77 6.45 5.95 6.10 7.77 11.46 11.63 7.09 10.34 34.00 55.76
$M_{\text{non}}(\%)\downarrow$ 2.13 2.33 2.50 3.63 1.51 2.76 3.16 3.33 3.63 3.88 3.04 2.91 2.00 3.75 10.83 14.62
Table 11: Cross-Country and Cross-Model Comparison of Win Percentages. This table presents the results for the Semantically Relevant, Culturally Relevant, and Non-Relevant Win Percentages across various multimodal retrievers for a selection of countries.
Appendix K Computational Resource
The experiment is performed with a single A100 GPU for approximately 3 gpu hours for each model or 54 hours in total with library version of colpali-engine 0.3.13.dev1+g9bee9b2b7, transformers 4.53.3 for most experiment except, GME models are inferenced under transformers 4.51.3
Appendix L Authoring and Implementation Tools
In preparing this manuscript, we utilized several generative large language models. For language editing and stylistic refinement, we employed Google’s Gemini 2.5-flash, along with models from xAI’s Grok family (e.g., Grok-3 Expert and Fast variants). For assistance with code implementation, scripting, and debugging, we used a model from Anthropic’s Claude series (e.g., Claude 4.0 Sonnet).
Appendix M Detailed Results

### v3

@@ Appendix C Evaluated Models and Details
   CLIP-L/14 clip-vit-large-patch141 427.6M
   CN-CLIP-L/14 Chinese-clip-vit-large-patch142 406.2M
@@
   XLM-R-L/14 XLM-Roberta-Large-Vit-L-146 998.3M
   XLM-R-L/16plus XLM-Roberta-Large-Vit-B-16Plus6 768.9M
@@
The models are based on the following works: 1) DBLP:journals/corr/abs-2103-00020 for CLIP-L/14; 2) chinese-clip for CN-CLIP-L/14; 3) faysse2025colpaliefficientdocumentretrieval for ColQwen2 models; 4) zhang2025gmeimprovinguniversalmultimodal for GME-Qwen2 models; 5) günther2025jinaembeddingsv4universalembeddingsmultimodal for Jina-E-v4; and 6) carlsson-etal-2022-cross for XLM-R-VL models.
@@
Appendix D Details on LBKL and DLBKL
In Section 3.1, we utilize LBKL and DLBKL to measure prevalence bias. Here, we elaborate on the rationale behind these metrics.
In an ideal, bias-free retrieval scenario, the distribution of languages in the retrieved list should match the expected ”fair” distribution (e.g., the ground truth distribution). LBKL measures the Kullback–Leibler divergence between these two distributions. In equation 1, $P(x)$ represents the proportion of a specific language in the ground truth, and $Q(x)$ represents the proportion in the retrieved set.
However, standard LBKL treats a deviation at rank 1 the same as a deviation at rank 100. In multimodal retrieval, users rarely scroll far, making top-rank bias significantly more harmful. Our proposed DLBKL in equation 2 addresses this by applying a discount factor derived from NDCG [ndcg]. By weighting the observed proportion $Q(x)$ by the inverse log of the rank, we ensure that high-resource language dominance at the top of the list results in a significantly higher divergence score than if that same dominance occurred lower in the list.
Appendix E Formulation of Language-wise LBKL and DLBKL
In this section, we detail the calculation of two bias metrics: Language-wise LBKL (Country-LBKL) and Language-wise DLBKL (Country-DLBKL). Both metrics quantify the divergence between the retrieved country distribution and an ideal uniform distribution. The key distinction lies in how the predicted distribution is constructed: Country-LBKL treats the retrieved documents as an unweighted set, while Country-DLBKL applies rank-based discounting to prioritize higher-ranked results.
E.1 Ground Truth Distribution (Shared)
Let $C=\{c_{1},c_{2},...,c_{N}\}$ be the set of target countries, where $N=|C|$. We define the ground truth distribution $P$ as:
$P(c_{j})=\frac{1}{N},\quad\forall c_{j}\in C$ (6)
E.2 Country-LBKL (Unweighted)
Country-LBKL evaluates bias based on the frequency of countries in the top-$k$ retrieved documents $D=[d_{1},...,d_{k}]$, treating all rank positions equally.
E.2.1 Predicted Distribution (Unweighted)
The unweighted proportion $Q(c_{j})$ is calculated as:
$Q(c_{j})=\frac{1}{k}\sum_{i=1}^{k}\mathbb{1}_{c_{j}}(d_{i})$ (7)
where the indicator function $\mathbb{1}_{c_{j}}(d_{i})$ is:
$\mathbb{1}_{c_{j}}(d_{i})=\begin{cases}1&\text{if }d_{i}\text{ is from }c_{j}\\ 0&\text{otherwise}\end{cases}$ (8)
E.2.2 Score Calculation
The Country-LBKL score is the KL divergence between $P$ and $Q$. We simplify the substitution of $P(c_{j})$ as follows:
$\begin{split}\text{Country-LBKL}&=\sum_{c_{j}\in C}P(c_{j})\log\left(\frac{P(c_{j})}{Q(c_{j})}\right)\\ &=\frac{1}{N}\sum_{c_{j}\in C}\log\left(\frac{1}{N\cdot Q(c_{j})}\right)\end{split}$ (9)
E.3 Country-DLBKL (Rank-Weighted)
Country-DLBKL assigns higher importance to documents appearing earlier in the list using logarithmic discounting.
E.3.1 Rank-Based Weighting
For rank position $i$, the weight $w_{i}$ is:
$w_{i}=\frac{1}{\log_{2}(i+1)}$ (10)
The total weight normalization factor is:
$W_{total}=\sum_{i=1}^{k}w_{i}=\sum_{i=1}^{k}\frac{1}{\log_{2}(i+1)}$ (11)
E.3.2 Predicted Distribution (Weighted)
The weighted distribution $Q^{\prime}(c_{j})$ is calculated as:
$Q^{\prime}(c_{j})=\frac{\sum_{i=1}^{k}w_{i}\cdot\mathbb{1}_{c_{j}}(d_{i})}{W_{total}}$ (12)
E.3.3 Score Calculation
The Country-DLBKL score is the KL divergence between $P$ and the weighted distribution $Q^{\prime}$:
$\begin{split}\text{Country-DLBKL}&=\sum_{c_{j}\in C}P(c_{j})\log\left(\frac{P(c_{j})}{Q^{\prime}(c_{j})}\right)\\ &=\frac{1}{N}\sum_{c_{j}\in C}\log\left(\frac{1}{N\cdot Q^{\prime}(c_{j})}\right)\end{split}$ (13)
E.4 Implementation Note
For both metrics, a small smoothing value is applied to $Q(c_{j})$ and $Q^{\prime}(c_{j})$ in implementation to ensure numerical stability when a country is not represented in the retrieval results.
Appendix F Example of Prevalence Bias Evaluation
To further elaborate on the result of research question 1, we provide the example of retrieval from image to text from CLIP and M-CLIP in Figure 8. We also provide the results of LBKL and DLBKL scores at other ranks in Table 8.
Figure 8: Qualitative Comparison of Image-to-Text Retrieval in Crossmodal-3600. Two retrieval results share identical semantic proportions and LBKL scores. Despite this semantic equality, the DLBKL score differs, capturing the model’s implicit preference for specific linguistic groups in the ranking process.
Model @5 @25 @50 @99
[…]
Vision-Language Contrastive Models
CLIP-L/14 15.846 15.848 14.706 14.711 14.158 14.162 13.629 13.631
CN-CLIP-L/14 15.784 15.787 14.419 14.427 13.680 13.688 12.968 12.974
Cross-lingual Alignment Models
XLM-R-L/14 14.710 14.718 8.547 8.583 4.759 4.800 2.386 2.413
XLM-R-B/16plus 14.654 14.662 8.248 8.285 4.345 4.388 2.007 2.036
MLLM-Based Retrieval Embedders
ColQwen2.5-3B-M 15.010 15.016 11.231 11.253 9.326 9.347 7.839 7.846
ColQwen2.5-7B-M 14.947 14.954 10.585 10.610 8.316 8.339 6.750 6.755
ColQwen2.5-v0.2 15.340 15.346 12.176 12.199 10.561 10.586 9.315 9.331
GME-Qwen2-2B 15.144 15.151 10.457 10.494 7.664 7.710 5.322 5.368
GME-Qwen2-7B 15.035 15.042 9.766 9.805 6.503 6.552 4.024 4.068
Jina-E-v4 14.781 14.789 9.320 9.354 6.079 6.120 3.734 3.769
Appendix G Language and Rank Frequency Diagram
To illustrate bias in image-to-text retrieval, we present a visualization of language groups categorized by resource level as shown in Appendix B, showing both their overall retrieval frequency as shown in Figure 9 and their frequency distribution across ranks as shown in Figure 10.
Figure 9: A language frequency histogram of each language group (see Appendix B) for all models
Figure 10: A frequency of language group (see Appendix B) at each rank for all model
Appendix H Culturally Relevant Concept Identification
To identify image concepts unique to each country, we first employed the Gemini2 as a tool for generating culturally relevant suggestions prior to data collection. The prompt used in this process is shown in Figure 11.
list 100 concepts that unique and vary in these country
including China, India, Japan, saudi arabia, France,
German, Brazil, Kenya, Thailand, USA
like this
{"food":{"China":"Mala Xiang Guo", ..., "Thailand":
"Padthai", "USA":"hamburger"}, "costume":{"China":
"Hanfu", ..., "Thailand":"Sabai", "USA":"cowboy"}}
Appendix I Multi-Concept Detection in Images
Please classify the following image by assigning them to one or more of the following cultural categories: {category}. **Comprehensive Output Format (JSON):** output as JSON for example:
  "<INDEX>": ["<CATEGORY>", "<CATEGORY>", ...],
in the categories please order by priority (high to low).
Appendix J Annotator Guideline
Appendix K Dataset Statistics
@@ Appendix I Multi-Concept Detection in Images
The distribution of cultural concepts in the 3XCM dataset is shown in Table 10, with each concept being represented by approximately 85 images on average.
Country Samples
@@
Average 732.75
Table 9: Dataset Statistics of 3XCM per Country
@@
Appendix L 3XCM Dataset Benchmark Samples
This research provides an association evaluation benchmark and image metadata. Examples of the benchmark for RQ2, RQ3, and image metadata are shown in Figure 13, 14, and 15 respectively.
Figure 13: Benchmark examples from the 3XCM dataset featuring three retrieval candidates.
Figure 14: Benchmark examples from the 3XCM dataset featuring six retrieval candidates.
Figure 15: Metadata for image of 3XCM dataset benchmark
Appendix M UMAP Analysis for Self-Preference Cultural Bias
To visualize cultural bias, the UMAP projections of text and image embeddings of all models as shown in Figure 16 and 17. The text embeddings cluster strongly by language, a proximity that supersedes semantic content. Conversely, the image embeddings do not exhibit strong country-based clustering, suggesting lower cultural bias. While other models show a similar, albeit less severe, tendency for text embeddings to be more biased than image embeddings, this effect is diminished in modern models. The GME-Qwen2 and Jina-E-v4 models only cluster very low-resource languages (Swahili, Yoruba), and the XLM-R models demonstrate superior alignment, forming a single central cluster. This discrepancy challenges retrieval systems: a query’s text embedding is biased by its language, leading the system to favor images from the same cultural context over potentially more visually relevant content from others.
Figure 17: The UMAP visualizations of the caption embeddings (left) and image embeddings (right) from the CN-CLIP-L/14 model applied to our dataset.
Appendix N Correlation Analysis for Self-Preference Cultural Bias
We investigate how unimodal bias, which our UMAP analysis shows is more severe in the text modality as shown in Appendix M, impacts cross-modal retrieval. To quantify this, we use the Silhouette score and find that high scores in low-resource languages correlate with the self-preference cultural bias score (SP), as shown in Table 12. We confirm this relationship by calculating the Pearson correlation between SP and the Silhouette scores. For example, CLIP-L/14’s Text Silhouette score correlates strongly with SP (0.827), while its Image Silhouette correlation is only moderate (0.550), as shown in Figure 18. Across all tested models, the average correlations reveal that SP is predominantly driven by the text encoder, as shown in Table 11.
Model TSC ISC
CLIP-L/14 0.83 0.55
CN-CLIP-L/14 -0.26 0.58
XLM-R-VL-B/16 0.98 -0.27
XLM-R-VL-L/14 0.94 0.05
Jina-E-v4 0.86 0.16
GME-Qwen2-2B 0.80 0.09
GME-Qwen2-7B 0.64 0.07
Average 0.68 0.18
@@
US GB AU DE CN JP FR ES AR PT BR SA TH IN KE NG
CN-CLIP-L/14 SP $\downarrow$ 0.03 0.06 0.05 1.11 0.03 0.13 0.33 0.40 0.43 0.56 0.52 4.88 2.55 6.98 1.99 2.23
Appendix O Detailed Results